Give vertex and axis of symmetry for f(x) = x ^2 + 2
(f(x)= x squared plus two
Not sure if the vertex should be (0,2) OR (0,-2) axis is 0.
I solved it by using the formula -b / 2a and then I substituted using the equation a^2 + bx +c.
thanks!

Answer by oscargut(686)

f(x) = x ^2 + 2
then -b/2a = -0/2=0
V=(0,f(0))=(0,2)
Axis: x=0